OpenStudy (anonymous):

There are 23 coins in a bank. All the coins are dimes and quarters. The total value of the coins is $3.80. How many dimes are there? How many quarters? How many dimes are in the coin bank?

OpenStudy (anonymous):

38 dimes 13 dimes 23 dimes none of the above

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

Not sure how you would figure this out without answers, but with the answers we know 38 doesnt equal 23 13 * 10 + ( (23-13) * 25) = 1.30 + 2.50 = 3.80 ** 23*10 doesn't equal 2.30 and there are no quarters. Its 13 dimes

OpenStudy (anonymous):

thanks :)